OpenStudy (anonymous):

what is the volume of the sphere? round the answer to the nearest cubic unit. Diameter= 18 cm

OpenStudy (anonymous):

3·π·93=3053.62806

OpenStudy (anonymous):

so 3054?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thanks. can u help with a few more?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

find the volume of the square pyramid. length=42 ft height=29ft

OpenStudy (anonymous):

So you want to do \[\frac{ l • h }{ 2 } or \frac{ 42 • 29 }{ 2 }\]

OpenStudy (anonymous):

5,600 ^3?

OpenStudy (anonymous):

where did you get the ^3 from?

OpenStudy (anonymous):

it is just one of my answer choices and i got 5600.. so idk where the ^3 came from

OpenStudy (anonymous):

Is 609 an answer?

OpenStudy (anonymous):

no.. it's only 5600, 11,760, 17, 052, 11,774

OpenStudy (anonymous):

ok hold on then

OpenStudy (anonymous):

Ok I was using the wrong formula.... It is 17,052

OpenStudy (anonymous):

thanks so much! can u help with a few more?

OpenStudy (anonymous):

Ya

OpenStudy (anonymous):

determine whether the two figures are similar. if so, give the scale factor of the smaller figure to the larger figure.

OpenStudy (anonymous):

I can't cuz I can't log in

OpenStudy (anonymous):

is it not just a pic like the last one?

OpenStudy (anonymous):

No try like dragging the image to your desktop or something

OpenStudy (anonymous):

?

OpenStudy (anonymous):

okay. well i can't seem to do it so i will just tell you. the height of the smaller figure is 9.9 and the radius is 1.3 and the height for the bigger figure is 23.76 and the radius is 3.12

OpenStudy (anonymous):

I don't think that they are. Those are some funky numbers tho?

OpenStudy (anonymous):

idk that is just what it says..

OpenStudy (anonymous):

I don't think they are similar

OpenStudy (anonymous):

Are they cylinders?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what is the scale factor?

OpenStudy (anonymous):

There is not a scale factor because they are not similar

OpenStudy (anonymous):

okay so the answer is no?

OpenStudy (anonymous):

can you see this image?

OpenStudy (anonymous):

Yes. and no

OpenStudy (anonymous):

ugh. okay well the next question is sorta the same

OpenStudy (anonymous):

smaller figure= height= 4, width =3, length =8 larger figure=height= 8, width=18, length= 72

OpenStudy (anonymous):

@jackmullen55 ?

OpenStudy (anonymous):

Rectangles?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

They don't look similar because 8:4 = 2:1 and 18:3 = 6:1 which means they are not of similar ratios

OpenStudy (anonymous):

alright thanks

OpenStudy (anonymous):

Yup!

OpenStudy (anonymous):

find the volume of the square pyramid. length = 12 cm, width = 12 cm height =11 cm